Fermi National Laboratory
 

Fermilab Shutdown Report for April 21, 2006 - April 28, 2006

 

LINAC

April 21 - 28

  • Running to NTF with the I- Ion Source
  • Beam Current 36 ma
  • Ready to Run 400 MeV Beam to the Dump

April 28 – May 5

  • Characterized LRF3 and LRF4
  • Replaced the Power Amplifier in LRF3 (it was sparking)
  • Replaced 4616 Driver In LRF4

Remaining Work

  • Make LRF4 Power Amplifier Run
  • Finish Characterization of the Low Energy RF Systems.
  • Calibrate High Energy Linac RF Gradient and Phase Readbacks
  • Install Motor Drive for 10 MeV Collimator (parts have arrived)
  • Find and Repair Ground fault in LRF7 Modulator.
  • Rebuild LRF7 Driver

Booster

  • MI-8 Dump cable pull finished
    • Terminating cables (upstairs/downstairs)
    • Vacuum continues to improve
    • Kickers installed/connected – pulsing starts today
    • Cleaning up tunnel
  • 400 MeV/ORBMP
    • Vacuum pump down continues
    • PS testing
    • Polarity Testing
  • RF repair/maintenance continues
  • Sump discharge re-route done today
  • RF repair and maintenance continues

Schedule

  • MI-8 Dump on schedule – (tasks remaining)
    • Terminate and test supplies
      • Septa, V809 switches, Kickers, Beam Stop
  • ORBMP – Finish PS work/testing
  • 400 MeV PS/Polarity Testing
  • RF work – Connect up Anode & 480
  • Safety System checkout
  • CUB LCW pumps – only 1 pump on now

Pbar

On Schedule

Main Injector

  • Continued vacuum repairs in the MI8 line.
  • Continued vacuum repairs at 620.
  • Began Lambertson bake at 101.
  • Continued installation of large aperture quad at 521,522.
  • Began installation of the MI10 kicker cooling system.

MI Startup (DRAFT)

Power Supply Startup:

  • Clear bus and Hipot.
  • Main PS commissioning (New wide aperture quads).
  • Establish ramps, slow magnet warm up.
  • Beamline PS turn on and commissioning.


Beam Startup:

  • Establish 8 GeV beam. 6 Hours.
  • MI8 line lattice measurements. 2x4 Hours.
  • Establish 120 GeV. beam. Smooth orbits. 8 Hours.
  • Commission lattice with new wide aperture quads 16 Hours.
  • High field Orbit smoothing (quad moves).  16 Hours.
  • Tune Up for Operation.    8 Hours.
  • Extraction line tune up. 4 hours each (7).


 Startup Activities and Studies:
Coalescing.

  • Damper commissioning.
  • NuMI Multibatch slip stacking.
  • SY 120
  • 8 GeV collimators
  • Install and commission new BPM electronics.
  • Install and commission new BLM electronics.
  • Test software programs with Linux.
  • Lambertson aperture scans.

Recycler

Tevatron

TEL 1

  • Installation underway

Jobs Not Yet Started

  • Realign low beta quads at B0 week of May 1
  • HOPs cable pull
  • A3 conning tower replacement, TD will to do conning towers in A3 next week
  • Unroll 48 magnets around ring
  • Measure rolls 13 magnets around ring

SwitchYard

Work continues.

NuMI

  • Horn 2 Repair – Successfully Tested the cooling
  • Drained and Refilled Horn 1 and 2 RAW Systems
  • Changed Horn 2 DI Bottle and High Potted OK
  • Planning on Horn 2 Pulsing for Next Week
  • Clapper Valve Installed on Horn 1 Raw System
  • Target Chase Chiller Delivery Expected Mid Next Week
  • Drains in Target Hall Being Covered to Reduce Area Humidity
  • Electric Installed and Ready for New Desiccant Units
  • Prism Well Underway with Dehumidification Duct Work
  • New Motorized Damper Installed Above Door to DK Passageway
  • OTR Installed and Surveyed in the Pre-Target
  • New Pre-Target Drip Ceiling Installed
  • Concrete Coring Complete for Various Piping and Electric
  • Excavation and Concrete Work in Process at SR3
    • Conduit In and Concrete Pour Today
  • Reworking HV Cables & Connections at New Motorized 13.8 Switch

MiniBooNE

  • Horn:
    • Complete (it works).
    • Reconnect old horn by May 8.
  • RAW System:
    • Complete.
  • Alignment:
    • Week of May 1.

Collider Detector at Fermilab

On schedule

  • Plug closes this week

D-Zero (D0)

  • Layer 0
    • Complete evaluation of Layer 0 readout performance
      • All SVX4 chips have been successfully readout with the sensors biased
    • Connect spool pieces and leak check joints
    • Re-install south inner H disk
  • Level 1 Calorimeter Trigger
    • Delivery of lower skew LVDS cables complete
    • Continue installing LVDS cables
    • Test transmission from ADF to TAB
    • Continue firmware debugging/development/verification
  • Level 1 Central Track Trigger
    • Finish connecting outputs to L1CalTrk and L1Muon
    • Verify outputs to L1CalTrk and L1Muon
  • Luminosity Monitors
    • Re-install and re-cable

Mechanical

Work Continues

Cryogenics

Cryo Operations

  • Tevatron at LN2 temperature and LN2 usage is stable with vendor deliveries running only 2-3 truckloads per day due to the Tevatron being 1/4 at room temperature. Helium losses are slightly up at 18 kscf/day (1/2 of normal during TeV operations); some of the loss was replenished with A0 GHe Recovery from the test caves.
  • A1 and A3 are at room temperature and isolated with LOTO.
  • D1 and F4 are also at room temperature, have been reconnected by Cryo, and are under helium flow. D1 is ready for Division sign-off.
  • F4 cooldown will wait until TEL-1 is re-installed.
  • B4 and E2 were cooled down to LN2 temperature.
  • A3 has been completed by MSD, but spool piece conning towers will be retrofitted (5 out of 8 leak) by Tech Division during this shutdown. Cryo has completed epoxy of helium leaks on Tevatron power lead and snoop leak tested fine.
  • Pbar D60 Debuncher tanks have been reconnected to system and we are awaiting Division sign-off before cooling down to LN2 temperature.
  • ACNET application programs relating to FRIG operations continue to be tested on Linux consoles.
  • Nitrogen Reliquefier startup is delayed due to concern of “old” electrical cables being used for cross-connect between MSS and KRSS, especially due to high startup currents of 3000 hp TSI compressor.
  • MCR monitoring cryo operations

CHL Maintenance:

  • General facility maintenance is in progress, including testing of relief valves per FESHM.
  • CHL Helium Compr “A” cooling tower installation has been completed with only some minor work is left, such as fencing; system testing will be commencing.
  • CHL Helium Compr “D” is being overhauled.
  • Helium Coldbox-2 controls upgrade is in progress.
  • Mobile Helium Purifier is being derimed at room temperature and piping modified.

FRIG Maintenance:

  • Various FRIG bldg vacuum maintenance and instrumentation repairs (e.g. lost charge on VPTs) are in progress, as well as support work on houses at room temperature and other equipment as things fail even during the shutdown period.
  • Various Mycom compressor maintenance: leak hunting and general repairs, F0#3 motor starter replacement in progress.
  • Wet expanders overhauls have been completed for A1 through F4; one D-Zero Dry Expander was also completed with modifications. AP30 and CDF remain to be done.
  • Kautzky relief valves have been replaced from C1 through F4 (16/24). Pace is at one house per day though running into problems with galling nuts on clamp's T-bolts and bad compression fittings. More Clamp T-bolts are on order.
  • We have been finding some leaking quick-disconnect fittings on conning towers. We have none in stock and thus ordered units expected next week.

Facility Engineering Services

  • Zebra mussel cleaning, though waiting for approval, is scheduled to begin on May 8 and continue for a minimum of 16 days.
  • Proposed Treatment
    • EVAC Introduction Stage
    • Flushing Stage
  • Hydrants Opened

  • Water Tested

Instrumentation

Work continues.

RF

Work continues.

Controls

Work continues.

ES&H Upgrades

Safety System Upgrades

Keep discussing safety with your crews to keep our focus on getting our jobs done safely.

  • Booster 60% Complete
    • Booster Dump Move Configuration Changes
      • LG1 area work started
        • Cables pulled to into the tunnel
        • New logic chassis being installed
        • Enclosure status pages updated for the new configuration
      • Booster Staging Area gate was replaced by mechanical support
    • Documentation submitted for Approval to Operate
  • Main Injector 40% Complete
    • Lower Level Controlled Access Displays Replaced
      • MI-62 & MI-30 complete. 5 Buildings left.
    • Main Injector Coasting Beam System changes
      • Lam10 Critical Device Interface Installation complete
    • Stairwell cleaning continues
      • MI Ring cleaning is complete.
      • Looking at the 8 GeV line stairwells.
  • P-Bar Source 100% Complete
    • Safety System Power Supply & Junction Box Replacements
    • Interlock New BSC700 Beam Stop
  • Tevatron 98% Complete
    • F-Sector Radiation Monitor Card Replacement
  • MT6 Section 1 & 2 70% Complete
    • Enclosure Interlock System Upgrade
  • Safety System Testing 0% Complete
    • All areas need to be re-tested before startup
    • Plan is to change Booster door cores Thursday evening
    • Booster Electrical Safety System testing is tentatively scheduled
      for 6:00 pm May 4th
  • Transition to Operations
    • Pressure to complete work to resume operations.
    • It’s the people that make the machines run.
    • Pressures to operate are not a reason to cut corners when it comes to working safely.

EE Support

Work continues

Alignment

  • Tevatron
    • Completed As found E10-E15.
    • Scheduled CDF Low Beta Alignment
  • Main Injector
    • Scheduled Wide Aperture Quad
    • Scheduled Align barrier cavity at MI604
  • Recycler
    • Completed Cooling Kicker Tank
    • Completed Alignment suspect magnets
  • 8Gev
    • Scheduled To Complete Pipe and New Toroid
    • Scheduled BPM Alignment to continue
  • P-Bar
    • Completed Girder Check
    • Scheduled ceramic beam tube alignment
  • MiniBooNE
    • Scheduled May 1
  • NuMI
    • Completed Re-Survey Multiwire and OTR.
    • Scheduled Horn check for Thursday
  • Booster
    • Not Scheduled
  • 400Mev
    • Scheduled Tuesday
  • Add Ons
    • D0 - Make Level-Run vertical measurement of select Collision Hall monuments West Wall
    • Main Injector Align barrier cavity at MI604 location.
    • Fess - the Batavia Rd. Guardhouse project.
    • Fess - MI Ponds C&D project will replace existing manholes and pipe.
    • Switchyard - Check the alignment of elements in Switchyard manhole enclosure F3.
    • Scheduled CDF Low Beta Alignment

Power Pole Replacement

  • 0700-0730 Master Sub goes off and this affects everything except for: Main Injector the Village, Mi 12a, Mi12b, Mi13, and Feeder 46a (cryo feeder).
  • 0700-0830 Feeder 48 (TeV RF) will go off.
  • At 0730-0745 Pbar AP10, 30, and AP50 will be moved to master substation AP0 will not be affected at this time.
  • At 0800-0830 Kautz Rd Substation off on Monday
  • 0900-1500 Power will be off in the Tevatron at E0, E1, E2, and E3 to replace a transformer.
  • (Special Note) Feeder 48 RF annual maintenance is currently scheduled for the day after all these substation outages from 0830-1700. (Currently Tuesday). Note: approximately a week later Feeder 48 goes down again

Pole Work

Schedule

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/28/2006    email Fermilab